A particle moves in a horizontal circular path of radius 2 m with uniform angular acceleration. It is observed to make 2 revolutions in the first 4 seconds of motion and 4 revolutions in the next 4 seconds. Find:
a) the initial angular velocity of the particle giving your answer in rad/s
b) the angular acceleration of the particle, giving your answer in rad per second squared.
1
Expert's answer
2020-06-11T10:34:29-0400

Let, a particle moves in circular motion with radius "r=2m"

Now, at "t=4s" ,angle traversed is "\theta_1=2\times2\pi=4\pi\:rad"

and at "t=8s" , angle traversed is "\theta_2=(2+4)\times 2\pi=12\pi \: rad"

Consider, the constant angular acceleration "\alpha" .


We know that,

"\theta=\omega_0+\frac{1}{2}\alpha t^2"

Thus,

"4\pi=\omega_0+8\alpha\hspace{1cm}(1)\\
12\pi=\omega_0+32\alpha\hspace{1cm}(2)"

(a).

multiply by 4 in equation (1) and subtract equation (2) from (1) we get

"\omega_0=\frac{4\pi}{3}\:rad/s"

(b).

Subtract equation (1) from(2) we get,

"\alpha=\frac{\pi}{3}\:rad/s^2"